# UTrace: Poisoning Forensics for Private Collaborative Learning

### v1

@@ 2.3 Related Work
Poisoning Traceback. A recent line of work on poisoning traceback complements existing defenses. Our work focuses on a recent setting [88, 40, 70] that aims to identify the source of poisoning attacks. Early works [88, 40] consider the problem of identifying training samples causing a malicious misclassification, and [70] consider traceback in MPC. We discuss these works in more detail in Section 4.1. Other works [13, 52] address poisoning traceback in federated learning. Other forensic settings for poisoning include reconstructing backdoor triggers [24] and traceback for privacy violations [63].
@@ 3.1 Problem Statement
PPML scenario. Figure 1 illustrates the general setup for PPML we consider. A set of several model owners host an ML service within a domain. For example, model owners might host a network threat detection tool for identifying malicious network activity from traffic logs containing private user information, or a cancer detection tool whose inputs contain private patient data. Such settings are typically subject to strict regulations restricting data use and sharing [2, 3].
@@ 4.1 Adapting Prior Techniques
We highlight some limitations of unlearning-based traceback in Appendix A. We observe that the approximate unlearning objective used by PF and Camel differs significantly from true unlearning, which can be exploited by an attacker. We also show that, in an idealized learning setting, using true unlearning with models trained with empirical risk minimization fails catastrophically if poisons are duplicated across even two user datasets (see Figure 5).
@@
The remaining obstacles with respect to challenges from Section 3.3 are effectiveness at low poisoning rates and the absence of user data during traceback. Intuitively, at low poisoning rates, the poisons are required to be more effective: hence, we expect poisons to appear even more suspicious, as measured by gradient similarity. Moreover, a useful property of gradient-based influence scores is that they only depends on gradient inner products. If gradients are cached during training, they can be used during traceback without relying on the presence of training data.
@@ 4.2 Extending Gradients to Users
[…]
@@ 4.3 Effective Scoring Function
Top-$k$ selection. Our main insight is to identify the most influential samples within each user’s data and include them in the aggregated user score. As $\mathcal{I}_{\cos}^{\mathrm{s}}$ aggregates GAS scores across samples, we propose to identify first the top-k samples with highest GAS scores and then aggregate only scores for those samples. If we define:
@@ 4.5 Framework Details
Training Routine. We describe the full UTrace training in Algorithm 1, incorporating all improvements discussed in Sections 4.3 and 4.4. The training output consists of both the trained model parameters $\theta_{t}$ and the training record $R$. Each element in $R$ contains the learning rate $\eta_{t}$, intermediate model parameters $\theta_{t-1}$, the training data gradients $g_{t}$, and the projection matrix $G_{t}$ for some training iteration $t\in\mathcal{T}$.
@@
Traceback Routine. We describe the UTrace traceback procedure in Algorithm 2. Given a misclassification event $z_{\mathrm{atk}}$, traceback computes user responsibility scores and outputs a ranked list of users based on these scores.
@@ 4.6 Secure UTrace
The main overhead of realizing Algorithm 2 as $\Pi_{\mathsf{\text{UTrace}}}$ (c.f. Protocol 3) comes from the gradient computation, computing the cosine distance between the gradients, and finding the $\mathrm{top\text{-}}k$ samples. The gradient computation is model-dependent and similar to training, so in the following we focus on realizing latter two computations efficiently. For the cosine similarity, a challenge is fixed-point division and approximating the divisor’s square roots. Integer division and square root operations require significantly more communication rounds compared to other operations, increasing both latency and bandwidth usage. Prior work proposed protocols for integer division and square roots by relying on Goldschmidt and Raphson-Newton iterations [6]. However, we can rewrite the gradient magnitudes in the divisor as a single square root operation due to the commutativity of multiplication with the square root for positive numbers.
@@
We provide a full description of this extension in Algorithm 4 in Section B.2. We implement $h$ as the ranking induced by the un-normalized gradient scores [78]:
@@
Realizing Selection in MPC. A challenge with realizing the heuristic in MPC is that the computation must be data-oblivious. In particular, the top-$l$ samples that we select with the heuristic cannot influence the subsequent computation, such as array accesses based on $\tilde{I}_{i}$. A naïve transformation to oblivious computation would entail doing the computation for all entries in $D_{i}$ and combining the result with an indicator bit based on whether we want to select the entry or not, but this removes the efficiency benefits of the algorithm, i.e., having to compute a cosine score for the full input array.
@@ 5.1 Evaluation setup
Datasets and models. We summarize the datasets and models we use in our experiments. Detailed training configurations, including hyperparameters, can be found in Appendix C.
@@
Traceback parameters. We report influence metric hyperparameters for unlearning and our method in Table 6 in Section C.1. In general, we sample checkpoints at some uniform interval until reaching a maximum checkpoint set size. We choose unlearning parameters using reasonable fine-tuning settings for each learning task.
@@ 5.2 Existing Poisoning Attacks on Vision
We begin our evaluation with standard attacks against CIFAR-10 and Fashion datasets. For space, we present detailed results for only CIFAR-10 here, and defer detailed Fashion results to Section D.2. Our qualitative findings are similar between both datasets.
Attack setup. We evaluate our tool against four different poisoning attacks: two backdoor attacks (BadNets [38] and Sleeper Agent [91]), Subpopulation [49], and targeted (Witches’ Brew [32]). BadNets and Subpopulation attacks are dirty-label attacks: the attacker has control over data labeling and can insert mislabeled points. Witches’ Brew and Sleeper Agent are clean-label attacks which work by applying small perturbations to inputs without changing their labels. Clean-label attacks are generally much more difficult to detect because they blend into the clean data distribution. We give complete details on attack setup in Section C.2.
Results. We report the ranking metrics for attacks against CIFAR-10 / ResNet-18 in Table 1 and the classification metrics in Table 2. We also plot attack-conditioned ROC curves in Figure 2.
@@
Parameter ablations. We vary the value of $k$ for the top-$k$ metric for different attacks. We find that setting $k$ too small ($k<16$ for CIFAR-10) weakens performance as the poisoning samples might not be included in the user score. On the other hand, if $k$ is too large, the benign samples influence the score more than poisoned samples. The value of $k=32$ provides good results for several attacks.
We run UTrace with 20 owners to evaluate the impact of the number of users. On BadNets, our method’s effectiveness deteriorates slightly, while unlearning maintains strong ranking. On Subpopulation attacks, our method improves from the 10 user setting, reaching $\geq 0.6$ TPR with 8 poisoned owners. Conversely, Unlearning degrades to a TPR of 0. Full results are in Section D.1.
@@ 5.3 Attacks against Unlearning
Attack setup. Motivated by our analysis in Appendix A, we implement two poisoning adversaries that prove challenging for unlearning. The first “Noisy BadNets” attack adds large amounts of label noise to backdoor samples with the goal of inducing a nearly uniform posterior distribution in the victim classifier on attack samples.
@@
Results. We report the ranking metrics for these attacks against CIFAR-10 / ResNet-18 in Table 3 and the classification metrics in Table 4. We also plot ROC curves in Figure 3.
@@ 5.4 Text and Malware Experiments
We present the results for experiments against the SST-2 and Ember datasets in Section D.3 and summarize the findings here. First, we find that the inner-product heuristic does not select relevant poisoning samples in the text modality. Hence, we use full $\mathrm{top\text{-}}k$ selection for attacks against SST-2. We experiment with four types of backdoor attack against SST-2 and two backdoor attacks against Ember. In general, we find the same trend of degradation in approximate unlearning as the number of poisoned datasets increases. For example, under a “style” text backdoor [76] against SST-2, our method achieves a TPR of 81.25%, versus the unlearning TPR of 4.7%. A similar result follows for an explanation-guided backdoor attack against Ember [86], where our method achieves a TPR of 98.4% versus the unlearning TPR of 5.9%.
@@ 5.5 Secure UTrace
Experimental Setup. We run UTrace on AWS c5.9xlarge machines running Ubuntu 20.04, each equipped with 36 vCPUs of an Intel Xeon 3.6 Ghz processor and 72 GB of RAM. The machines are connected over a local area network (LAN) through a 12 Gbps network interface with an average round-trip time (RTT) of $0.5$ ms. We additionally perform our experiments in a simulated wide area network (WAN) setting using the tc utility to introduce an RTT of $80$ ms and limit the bandwidth to $2$ Gbps. We report the total wall clock time and the communication cost based on the data sent by each party, which includes the time and bandwidth required for setting up any cryptographic material for the MPC protocol, such as correlated randomness. As both our algorithm and Camel contain parts that can be reused across computations, we split the cost into preprocessing and online phases.
@@
The wall-clock time and bandwidth results shown in Figure 4(a) highlight this improvement. Roughly speaking, for Camel, the additional $m\cdot e$ training epochs is similar to performing a full training run, taking 696 hours and 2734 hours for semi-honest and malicious protocols, respectively, for CIFAR-10 in the LAN setting. UTrace’s overhead shows an $2-3.5$x reduction over Camel’s, taking only 374 hours and 1372 hours for semi-honest and malicious protocols, respectively. In conclusion, the cost of the traceback procedure is only a fraction of the MPC training cost.
Online. We compare the wall-clock time and bandwidth costs of the online time of the strawman, the optimization and the heuristic versions of UTrace, as well as Camel in Figure 4(b). The overhead of Camel is 27 seconds for the malicious protocol in LAN and 73 minutes in WAN for CIFAR-10, because it only requires $O(m)$ forward passes of the test sample, i.e., one inference per input party. UTrace, on the other hand, requires $O(|\mathcal{T}|)$ forward passes in order to compute the gradients for each checkpoint, which take 137 seconds for the malicious protocol in LAN and 350 minutes in WAN. The reason for Camel’s faster latency is that the number of parties $m$ is smaller than the number of checkpoints $|\mathcal{T}|$ in our setting.
In addition to the inference, computing cosine distances introduces an additional overhead of $1.8-4.2$x (an extra $249-314$ seconds) in the malicious LAN setting. We reduce this overhead to $0.58-1.41$x in LAN using the reciprocal square root optimization. UTrace with the heuristic optimization further reduces this overhead to $0.36-0.94$x of the forward pass cost. In absolute terms, the cost of doing a traceback is 44-48 seconds (105-106 minutes) and 187-217 seconds (370-374 minutes), respectively, for the semi-honest and malicious protocols in the LAN (WAN) setting for CIFAR-10. Considering the cost of a single inference in the LAN (WAN) setting, 0.83 and 2.25 seconds (2 and 6.8 minutes), the overhead of running UTrace is reasonable. Note that these numbers include the cost to compute input-independent cryptographic material of the underlying MPC protocols, and MPC-protocol-dependent preprocessing optimizations can be applied to reduce latency further.
@@ 6 Conclusions, Limitations, and Future Work
We design a user-level traceback procedure UTrace in PPML with high effectiveness and low FPR that generalizes across backdoor, targeted, and subpopulation attacks. Compared to unlearning-based methods such as Camel [70], UTrace achieves higher TPR at low FPR, and provides a better ranking of malicious users. Our optimizations make the tool practical in PPML settings, as the pre-processing cost is $2-3.5$x lower than that of Camel, at the cost of a slight increase in online traceback cost.
@@ Acknowledgments
This research was sponsored by NSF awards CNS-2312875 and CNS-23121720, and the contract W911NF2420115, issued by US ARMY ACC-APG-RTP.
@@ Appendix A Analysis of Unlearning
In Section A.1, We examine the basic assumptions made by approximate unlearning, and discuss scenarios where they fail. We pair these findings with experimental results demonstrating the failure modes with concrete attacks. These results demonstrate that there exist attacks, some natural and some adaptive, which evade detection by approximate unlearning, motivating the need for better responsibility scores.
In Section A.2, we state and prove a theorem characterizing a failure mode of true unlearning for user-level detection; that is, in the case when the approximate unlearning routine $\operatorname{Unl}(\theta;D^{\prime};D_{\mathrm{tr}})$ is replaced with retraining from-scratch on the dataset $D^{\prime}\setminus D_{\mathrm{tr}}$. This result implies that the vulnerabilities of approximate unlearning cannot be resolved simply by resorting to a more accurate unlearning procedure. In fact, approximate unlearning provides resilience to multiple poisoned owners.
@@ A.1 Approximate Unlearning Failure Modes
Let $\mu_{1},\ldots,\mu_{n}$ be distributions over $\mathcal{Z}$, and let $10<\alpha_{1},\ldots,\alpha_{n}<$ 1 be mixing weights (i.e., $\sum_{i}\alpha_{i}=1$). Let $\nu=\sum_{i}\alpha_{i}\mu_{i}$. Then, for all $(x,y)\in\mathcal{Z}$,
@@
Lemma A.1 is a straightforward consequence of Bayes’ Theorem, and describes the posterior $\nu(y|x)$ as a convex combination of posteriors $\mu_{1}(y|x),\ldots,\mu_{n}(y|x)$. In particular, the posteriors under both $\nu^{\mathrm{unl}}_{p}$ and $\nu^{\mathrm{unl}}_{c}$ are given by
@@
where $C$ is the number of classes. We can inspect Equations 11 and 12 for certain representative poisoning settings. For instance, for classification tasks with unambiguous labeling, it is likely that $\mu_{c}(y^{\mathrm{atk}}|x^{\mathrm{atk}})=0$, and so the posteriors simplify to

### v2

# procbcenter,boxedcodesize=,width=

@@ 2.3 Related Work
Poisoning Traceback. A recent line of work on poisoning traceback complements existing defenses. Our work focuses on a recent setting [88, 40, 70] that aims to identify the source of poisoning attacks. Early works [88, 40] consider the problem of identifying training samples causing a malicious misclassification, and [70] consider traceback in MPC. We discuss these works in more detail in Section˜4.1. Other works [13, 52] address poisoning traceback in federated learning. Other forensic settings for poisoning include reconstructing backdoor triggers [24] and traceback for privacy violations [63].
@@ 3.1 Problem Statement
PPML scenario. Figure˜1 illustrates the general setup for PPML we consider. A set of several model owners host an ML service within a domain. For example, model owners might host a network threat detection tool for identifying malicious network activity from traffic logs containing private user information, or a cancer detection tool whose inputs contain private patient data. Such settings are typically subject to strict regulations restricting data use and sharing [2, 3].
@@ 4.1 Adapting Prior Techniques
We highlight some limitations of unlearning-based traceback in Appendix˜A. We observe that the approximate unlearning objective used by PF and Camel differs significantly from true unlearning, which can be exploited by an attacker. We also show that, in an idealized learning setting, using true unlearning with models trained with empirical risk minimization fails catastrophically if poisons are duplicated across even two user datasets (see Figure˜5).
@@
The remaining obstacles with respect to challenges from Section˜3.3 are effectiveness at low poisoning rates and the absence of user data during traceback. Intuitively, at low poisoning rates, the poisons are required to be more effective: hence, we expect poisons to appear even more suspicious, as measured by gradient similarity. Moreover, a useful property of gradient-based influence scores is that they only depends on gradient inner products. If gradients are cached during training, they can be used during traceback without relying on the presence of training data.
@@ 4.2 Extending Gradients to Users
Building on our insights, gradient-based influence scores appear the most promising for implementing poisoning traceback in PPML. The main challenge is how to adapt gradient similarity techniques to operate at the user level. We begin by recalling the GAS score [40]. Before training, an iteration set $\mathcal{T}\subseteq[T]$ is fixed, where $T$ is the number of training steps. Typically $|\mathcal{T}|\ll T$. During training, at iterations $t\in\mathcal{T}$, the intermediate model parameters $\theta_{t}$ are recorded along with the current learning rate $\eta_{t}$. The sample-level influence score $\mathcal{I}_{\cos}$ between training point $z$ and test point $\hat{z}$ is then defined as
@@ 4.3 Effective Scoring Function
[…]
@@ 4.5 Framework Details
Training Routine. We describe the full UTrace training in Algorithm˜1, incorporating all improvements discussed in Sections˜4.3 and 4.4. The training output consists of both the trained model parameters $\theta_{t}$ and the training record $R$. Each element in $R$ contains the learning rate $\eta_{t}$, intermediate model parameters $\theta_{t-1}$, the training data gradients $g_{t}$, and the projection matrix $G_{t}$ for some training iteration $t\in\mathcal{T}$.
@@
Traceback Routine. We describe the UTrace traceback procedure in Algorithm˜2. Given a misclassification event $z_{\mathrm{atk}}$, traceback computes user responsibility scores and outputs a ranked list of users based on these scores.
@@ 4.6 Secure UTrace
The main overhead of realizing Algorithm˜2 as $\Pi_{\mathsf{\text{UTrace}}}$ (c.f. Protocol 3) comes from the gradient computation, computing the cosine distance between the gradients, and finding the $\mathrm{top\text{-}}k$ samples. The gradient computation is model-dependent and similar to training, so in the following we focus on realizing latter two computations efficiently. For the cosine similarity, a challenge is fixed-point division and approximating the divisor’s square roots. Integer division and square root operations require significantly more communication rounds compared to other operations, increasing both latency and bandwidth usage. Prior work proposed protocols for integer division and square roots by relying on Goldschmidt and Raphson-Newton iterations [6]. However, we can rewrite the gradient magnitudes in the divisor as a single square root operation due to the commutativity of multiplication with the square root for positive numbers.
@@
We provide a full description of this extension in Algorithm˜4 in Section˜B.2. We implement $h$ as the ranking induced by the un-normalized gradient scores [78]:
@@
[…]
@@ 5.1 Evaluation setup
Datasets and models. We summarize the datasets and models we use in our experiments. Detailed training configurations, including hyperparameters, can be found in Appendix˜C.
@@
Traceback parameters. We report influence metric hyperparameters for unlearning and our method in Table˜6 in Section˜C.1. In general, we sample checkpoints at some uniform interval until reaching a maximum checkpoint set size. We choose unlearning parameters using reasonable fine-tuning settings for each learning task.
@@ 5.2 Existing Poisoning Attacks on Vision
We begin our evaluation with standard attacks against CIFAR-10 and Fashion datasets. For space, we present detailed results for only CIFAR-10 here, and defer detailed Fashion results to Section˜D.2. Our qualitative findings are similar between both datasets.
Attack setup. We evaluate our tool against four different poisoning attacks: two backdoor attacks (BadNets [38] and Sleeper Agent [91]), Subpopulation [49], and targeted (Witches’ Brew [32]). BadNets and Subpopulation attacks are dirty-label attacks: the attacker has control over data labeling and can insert mislabeled points. Witches’ Brew and Sleeper Agent are clean-label attacks which work by applying small perturbations to inputs without changing their labels. Clean-label attacks are generally much more difficult to detect because they blend into the clean data distribution. We give complete details on attack setup in Section˜C.2.
Results. We report the ranking metrics for attacks against CIFAR-10 / ResNet-18 in Table˜1 and the classification metrics in Table˜2. We also plot attack-conditioned ROC curves in Figure˜2.
@@
[…]
We run UTrace with 20 owners to evaluate the impact of the number of users. On BadNets, our method’s effectiveness deteriorates slightly, while unlearning maintains strong ranking. On Subpopulation attacks, our method improves from the 10 user setting, reaching 0.6$\geq$ 0.6≥ 0.6 TPR with 8 poisoned owners. Conversely, Unlearning degrades to a TPR of 0. Full results are in Section˜D.1.
@@ 5.3 Attacks against Unlearning
Attack setup. Motivated by our analysis in Appendix˜A, we implement two poisoning adversaries that prove challenging for unlearning. The first “Noisy BadNets” attack adds large amounts of label noise to backdoor samples with the goal of inducing a nearly uniform posterior distribution in the victim classifier on attack samples.
@@
Results. We report the ranking metrics for these attacks against CIFAR-10 / ResNet-18 in Table˜3 and the classification metrics in Table˜4. We also plot ROC curves in Figure˜3.
@@ 5.4 Text and Malware Experiments
We present the results for experiments against the SST-2 and Ember datasets in Section˜D.3 and summarize the findings here. First, we find that the inner-product heuristic does not select relevant poisoning samples in the text modality. Hence, we use full $\mathrm{top\text{-}}k$ selection for attacks against SST-2. We experiment with four types of backdoor attack against SST-2 and two backdoor attacks against Ember. In general, we find the same trend of degradation in approximate unlearning as the number of poisoned datasets increases. For example, under a “style” text backdoor [76] against SST-2, our method achieves a TPR of 81.25%, versus the unlearning TPR of 4.7%. A similar result follows for an explanation-guided backdoor attack against Ember [86], where our method achieves a TPR of 98.4% versus the unlearning TPR of 5.9%.
@@ 5.5 Secure UTrace
Experimental Setup. We run UTrace on AWS c5.9xlarge machines running Ubuntu 20.04, each equipped with 36 vCPUs of an Intel Xeon 3.6 Ghz processor and 72 GB of RAM. The machines are connected over a local area network (LAN) through a 12 Gbps network interface with an average round-trip time (RTT) of 0.50.50.5 ms. We additionally perform our experiments in a simulated wide area network (WAN) setting using the tc utility to introduce an RTT of 808080 ms and limit the bandwidth to 222 Gbps. We report the total wall clock time and the communication cost based on the data sent by each party, which includes the time and bandwidth required for setting up any cryptographic material for the MPC protocol, such as correlated randomness. As both our algorithm and Camel contain parts that can be reused across computations, we split the cost into preprocessing and online phases.
@@
The wall-clock time and bandwidth results shown in Figure˜4(a) highlight this improvement. Roughly speaking, for Camel, the additional $m\cdot e$ training epochs is similar to performing a full training run, taking 696 hours and 2734 hours for semi-honest and malicious protocols, respectively, for CIFAR-10 in the LAN setting. UTrace’s overhead shows an 23.52-3.52 - 3.5x reduction over Camel’s, taking only 374 hours and 1372 hours for semi-honest and malicious protocols, respectively. In conclusion, the cost of the traceback procedure is only a fraction of the MPC training cost.
Online. We compare the wall-clock time and bandwidth costs of the online time of the strawman, the optimization and the heuristic versions of UTrace, as well as Camel in Figure˜4(b). The overhead of Camel is 27 seconds for the malicious protocol in LAN and 73 minutes in WAN for CIFAR-10, because it only requires $O(m)$ forward passes of the test sample, i.e., one inference per input party. UTrace, on the other hand, requires $O(|\mathcal{T}|)$ forward passes in order to compute the gradients for each checkpoint, which take 137 seconds for the malicious protocol in LAN and 350 minutes in WAN. The reason for Camel’s faster latency is that the number of parties $m$ is smaller than the number of checkpoints $|\mathcal{T}|$ in our setting.
In addition to the inference, computing cosine distances introduces an additional overhead of 1.84.21.8-4.21.8 - 4.2x (an extra 249314249-314249 - 314 seconds) in the malicious LAN setting. We reduce this overhead to 0.581.410.58-1.410.58 - 1.41x in LAN using the reciprocal square root optimization. UTrace with the heuristic optimization further reduces this overhead to 0.360.940.36-0.940.36 - 0.94x of the forward pass cost. In absolute terms, the cost of doing a traceback is 44-48 seconds (105-106 minutes) and 187-217 seconds (370-374 minutes), respectively, for the semi-honest and malicious protocols in the LAN (WAN) setting for CIFAR-10. Considering the cost of a single inference in the LAN (WAN) setting, 0.83 and 2.25 seconds (2 and 6.8 minutes), the overhead of running UTrace is reasonable. Note that these numbers include the cost to compute input-independent cryptographic material of the underlying MPC protocols, and MPC-protocol-dependent preprocessing optimizations can be applied to reduce latency further.
@@ 6 Conclusions, Limitations, and Future Work
We design a user-level traceback procedure UTrace in PPML with high effectiveness and low FPR that generalizes across backdoor, targeted, and subpopulation attacks. Compared to unlearning-based methods such as Camel [70], UTrace achieves higher TPR at low FPR, and provides a better ranking of malicious users. Our optimizations make the tool practical in PPML settings, as the pre-processing cost is 23.52-3.52 - 3.5x lower than that of Camel, at the cost of a slight increase in online traceback cost.
@@ Acknowledgments
This research was sponsored by NSF awards CNS-2312875 and CNS-2331081, and the contract W911NF2420115, issued by US ARMY ACC-APG-RTP.
@@ Appendix A Analysis of Unlearning
In Section˜A.1, We examine the basic assumptions made by approximate unlearning, and discuss scenarios where they fail. We pair these findings with experimental results demonstrating the failure modes with concrete attacks. These results demonstrate that there exist attacks, some natural and some adaptive, which evade detection by approximate unlearning, motivating the need for better responsibility scores.
In Section˜A.2, we state and prove a theorem characterizing a failure mode of true unlearning for user-level detection; that is, in the case when the approximate unlearning routine $\operatorname{Unl}(\theta;D^{\prime};D_{\mathrm{tr}})$ is replaced with retraining from-scratch on the dataset $D^{\prime}\setminus D_{\mathrm{tr}}$. This result implies that the vulnerabilities of approximate unlearning cannot be resolved simply by resorting to a more accurate unlearning procedure. In fact, approximate unlearning provides resilience to multiple poisoned owners.
@@ A.1 Approximate Unlearning Failure Modes
Let $\mu_{1},\ldots,\mu_{n}$ be distributions over $\mathcal{Z}$, and let $0<\alpha_{1},\ldots,\alpha_{n}<1$ be mixing weights (i.e., $\sum_{i}\alpha_{i}=1$). Let $\nu=\sum_{i}\alpha_{i}\mu_{i}$. Then, for all $(x,y)\in\mathcal{Z}$,
@@
Lemma˜A.1 is a straightforward consequence of Bayes’ Theorem, and describes the posterior $\nu(y|x)$ as a convex combination of posteriors $\mu_{1}(y|x),\ldots,\mu_{n}(y|x)$. In particular, the posteriors under both $\nu^{\mathrm{unl}}_{p}$ and $\nu^{\mathrm{unl}}_{c}$ are given by
@@
where $C$ is the number of classes. We can inspect Equations˜11 and 12 for certain representative poisoning settings. For instance, for classification tasks with unambiguous labeling, it is likely that $\mu_{c}(y^{\mathrm{atk}}|x^{\mathrm{atk}})=0$, and so the posteriors simplify to